If I have 117 Quaters and dimes together and the amount of money together is $15.75 how many of each coin do I have
Answer by jorel555(1290)

Let q be quarters, and d be dimes. Then:
d+q=117
d=117-q
10(117-q)+25q=1575
1170+15q=1575
15q=405
q=27
d=90
You have 90 dimes and 27 quarters totaling $15.75. ☺☺☺☺